OpenStudy (username123):

SOMEBODY ANYBODY EVERYBODY HELP!! Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7.

OpenStudy (username123):

y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (username123):

i have no idea how to do this

OpenStudy (username123):

which answer is it tho

OpenStudy (anonymous):

plug in your given point and see which one line is satisfied

OpenStudy (anonymous):

You know the slope should be 1, so use point-slope formula. (y-y1)=m(x-x1) y-(-2)=1(x-6) now solve for y

OpenStudy (username123):

is it y=x-8

OpenStudy (anonymous):

is that what you got when you solved for y? if so, then yes.

OpenStudy (username123):

yes! Thanks!